KyledaGreat:

Consider the following relation. x=−4|y|+3 Find four points contained in the inverse. Express your values as an integer or simplified fraction.

KyledaGreat:

@tranquility

Tranquility:

This question looks familiar. What is the inverse?

KyledaGreat:

-4x + 3

Tranquility:

Don't forget to keep the y = Also there should be absolute values around the x y = -4|x| + 3

KyledaGreat:

x values: -2, -1, 0, 1, 2

Tranquility:

Do you know how absolute value works? When you're plugging in those numbers to find the corresponding y-values: y = -4 |x| + 3 When x is -2 y = -4 |(-2)| + 3 The absolute value of -2 is going to be 2. Any negative number becomes positive Positive numbers stay the same y = -4*2 + 3 y = -8 + 3 y = -5 The point is (-2, -5)

Tranquility:

Repeat this for x = -1, 0, and 1

KyledaGreat:

When x is -1 y = -4*1 + 3 y = 7 + 3 y = 10 (-1, 10)

Tranquility:

Check your work again -4* 1 is -4 What is -4 + 3

KyledaGreat:

-1

Tranquility:

So your point is going to be (-1, -1) You can also check the graph: https://www.desmos.com/calculator/uvmvin3tiy

KyledaGreat:

When x is 0 y = -4*0+ 3 y = 3 + 3 y = 6 (0, 6)

Tranquility:

Where did you get 3 + 3 from?

Tranquility:

y = (-4*0)+ 3 Simplify it again

KyledaGreat:

3

Tranquility:

Now you got it. What's the point if y is 3 when x is 0

KyledaGreat:

i'm not sure to be honest

KyledaGreat:

3 + 0

KyledaGreat:

is this right

Tranquility:

(x, y) When x is 0, y is 3 (0, 3)

KyledaGreat:

1 When x is 1 y = -4*1+ 3 y = 1 + 3 y = 4

Tranquility:

Once again, I'm not sure why you've been making the same mistake again and again

Tranquility:

y = -4|x| + 3 when x = 1 y = -4|1| + 3 |1| is equal to 1 so y = (-4*1) + 3 what is y equal to?

KyledaGreat:

-1

Tranquility:

That is correct

Tranquility:

Just think of the | | as some sort of special parenthesis which just turns the number inside to become positive

KyledaGreat:

oh okay

KyledaGreat:

(1,-1)

Tranquility:

Yes

KyledaGreat:

Consider the following relation. x=−4|y|+3 Find the domain and range of the inverse. Express your answer in interval notation.

KyledaGreat:

Domain : \[(-\infty,\infty)\]

KyledaGreat:

i don't think so if the range is right \[(-\infty, 3]\]

KyledaGreat:

is that right

KyledaGreat:

@tranquility

Tranquility:

@kyledagreat wrote:
Domain : \[(-\infty,\infty)\]
@kyledagreat wrote:
i don't think so if the range is right \[(-\infty, 3]\]
That is correct